The dean of a college is interested in the proportion of graduates from his college who have a job offer on graduation day. He is particularly interested in seeing if there is a difference in this proportion for accounting and economics majors. In a random sample of 200 of each type of major at graduation, he found that 72 accounting majors and 52 economics majors had job offers. Assume pooled estimate of the population proportion and a level of significance (α) of 0.01.

1. State your null and alternate hypotheses:

2. What is the value of the test statistic? Please show all the relevant calculations.

3. What is the rejection criterion based on the p-value approach?

4. What is the Statistical decision (i.e. reject /or do not reject the null hypothesis) based on the p -value obtained?
Answer #1

let p1 and p2 are proportion of accounting and economic majors that have job offers

null Hypothesis alternate Hypothesis: Ho: P1-P2 0.0000 Ha P1-P2 0.0000 for 0.01 level with two tailed test, critical value of z= Decision rule 2.576 reject Ho if absolute value of test statistic 리 >2.576 Accounting 72 0.3600 200 Economic 52 0.2600 200 0.1000 0.3100 estimated difference in proportion estimator of pooled proportion p =(X1+X2)/(n-m2) p,-Pz std error Se= 0.0462 test statz- 2.16 P value 0.0308

from above

1)

null Hypothesis:               Ho:    p1-p2 = 0.0000
alternate Hypothesis:    Ha:    p1-p2 0.0000

2)

test statistic z=2.16

3) rejection criterion : if p value <0.01 level ; reject Ho

4)

as p value(0.0308) is not less then significance level 0.01 ; we can not reject null hypothesis